### v1

# Domain Generalization in-the-Wild: Disentangling Classification from Domain-Aware Representations

Abstract
Evaluating domain generalization (DG) for foundational models like CLIP is challenging, as web-scale pretraining data potentially covers many existing benchmarks. Consequently, current DG evaluation may neither be sufficiently challenging nor adequately test genuinely unseen data scenarios. To better assess the performance of CLIP on DG in-the-wild, a scenario where CLIP encounters challenging unseen data, we consider two approaches: (1) evaluating on 33 diverse datasets with quantified out-of-distribution (OOD) scores after fine-tuning CLIP on ImageNet, and (2) using unlearning to make CLIP ‘forget’ some domains as an approximation. We observe that CLIP’s performance deteriorates significantly on more OOD datasets. To address this, we present CLIP-DCA (Disentangling Classification from enhanced domain Aware representations). Our approach is motivated by the observation that while standard domain invariance losses aim to make representations domain-invariant, this can be harmful to foundation models by forcing the discarding of domain-aware representations beneficial for generalization. We instead hypothesize that enhancing domain awareness is a prerequisite for effective domain-invariant classification in foundation models. CLIP-DCA identifies and enhances domain awareness within CLIP’s encoders using a separate domain head and synthetically generated diverse domain data. Simultaneously, it encourages domain-invariant classification through disentanglement from the domain features. CLIP-DCA shows significant improvements within this challenging evaluation compared to existing methods, particularly on datasets that are more OOD.
1 Introduction
Domain generalization (DG) aims to train models that maintain robust performance when encountering out-of-distribution (OOD) data zhou2022domain . A key assumption of DG is that the target domains represent novel data distributions for evaluation. However, this assumption is challenged when evaluating pretrained foundation models like CLIP radford2021learning_CLIP and ALIGN jia2021scaling . These models have been trained on comprehensive web-scale datasets, thus have likely been exposed to most existing domains, contributing to its impressive zero-shot capabilities. Consequently, much research has focused on adapting CLIP through parameter-efficient finetuning zhou2022learning_coop ; zhou2022conditional_cocoop ; gao2024clipadapter ; zhang2022tip , regularization using the original weights wortsman2022robust_wise ; nam2024lipsum ; oh2024towards_carot ; shu2023clipood , and even transductive methods wallingford2023neuralpriming ; martin2024transductive , largely preserving its pretrained knowledge. However, a critical question arises: how would CLIP perform on genuinely unseen domains? A recent study mayilvahanan2024search found that retraining CLIP from scratch using only natural images significantly degrades performance on OOD benchmarks, resulting in performance similar to models trained only on ImageNet. Our results (Figure 1) align with these findings, suggesting current DG evaluations for CLIP may overestimate its true OOD robustness because standard evaluation settings like leave-one-domain-out and existing cross-dataset evaluations may not be sufficiently challenging (Sec. 4.1, 4.2).
Figure 1: Improvement over zeroshot after finetuning on ImageNet (in %). OOD scores are quantified relative to ImageNet (source dataset), illustrating the challenge of DG in-the-wild.
We therefore propose that DG evaluation for foundation models, such as CLIP, should be more challenging, to approximate “domain generalization in-the-wild," where CLIP might encounter diverse and challenging new data in the real-world. We evaluate CLIP on 33 target datasets spanning a diverse range of OODness. To systematically approach evaluation, we quantify a multi-modal OOD score (Sec. 4.2), using ImageNet as both an anchor and a source dataset owing to its inclusion of many classes and concepts. We find that after finetuning on ImageNet, CLIP’s DG performance degrades on datasets with higher OOD scores with respect to ImageNet (Figure 1), consistent with the domain contamination findings mayilvahanan2024search . In addition, to further simulate truly unseen domains, we use an unlearning technique sepahvandselective to make CLIP forget some domains (Sec. 4.3), and find significant performance degradation for existing robust finetuning methods.
Our results (Figure 9), alongside findings on domain contamination mayilvahanan2024search , suggest that for DG in-the-wild, different robust finetuning algorithms are needed for genuinely unseen data. In light of this, we present CLIP-DCA (Disentangling Classification from enhanced domain Aware representations), an end-to-end finetuning method to improve the robustness of CLIP on truly OOD data. A key idea in DG is that learning domain-invariant features is beneficial for robust generalization zhou2022domain ; ganin2016domain_dann . However, naively enforcing domain invariance for a pretrained foundation model could cause catastrophic forgetting of useful features learned from diverse domains during pretraining as the model is forced to make its representations entirely domain-invariant. We hypothesize that to learn effective domain invariance, domain awareness is a prerequisite. This awareness is critical to maintain CLIP’s vast knowledge, which includes generalizable features that support capabilities like zero-shot classification. By enhancing domain awareness, CLIP can also selectively disentangle classification from domain-specific aspects, thereby achieving robust generalization without forgetting valuable information.
We combine the idea of domain awareness and domain invariance by encouraging them simultaneously within CLIP-DCA (Figure 2). Specifically, we encourage domain awareness within CLIP’s image and text encoders, while promoting domain invariance specifically at the final classification layer through disentanglement. Our premise is that while domain awareness is a requirement to maintain pre-existing knowledge, this awareness can be disentangled for domain-invariant classification and robust generalization. To achieve this, we add a new head to the CLIP image encoder, called the domain head, which is trained to understand domains. The original classification head is then disentangled from the domain head, effectively learning domain awareness within its encoders and achieving domain invariance at the classification stage. Additionally, since many datasets lack distinct domains or textual descriptions, and the definition of ‘domain’ is often vague in DG in-the-wild, we address this by using diffusion models to create images of artificial domains and MLLMs to generate descriptions for these artificial domains (Sec. 3.2). Our contributions are summarized as follows:
  • We demonstrate potential limitations in current DG evaluations of foundation models, supported by our results and a recent study. Existing benchmarks may overestimate true OOD robustness, potentially leading finetuning strategies towards in-distribution improvement rather than OOD.
  • We propose more challenging and holistic evaluations for DG in-the-wild. We use an expanded cross-dataset evaluation setting spanning 33 datasets from diverse domains, indexed by multi-modal OOD scores. We also use an unlearned model to further approximate unseen domains.
  • We introduce CLIP-DCA, a novel finetuning method that improves OOD robustness by disentangling classification from enhanced domain-aware representations. We find that on more OOD target datasets, CLIP-DCA performs significantly better compared to existing robust finetuning methods, while performance is similar across all methods on less OOD target datasets.
2 Related Work
Domain Generalization. Learning domain-invariant representations has historically been a central idea in domain generalization blanchard2011generalizing ; zhou2022domain . The intuition is that when classifying images from entirely new distributions, learning abstract features common across source domains should provide better robustness for classification in new domains blanchard2011generalizing ; muandet2013domain . Among these, domain-adversarial learning methods have become a relatively standard approach within the DG field due to its conceptual simplicity and effectiveness zhou2022domain . For instance, Domain Adversarial Neural Networks (DANN) ganin2016domain_dann uses an auxiliary domain classifier trained adversarially against the encoder, encouraging the encoder to produce features indistinguishable across source domains. Given the focus of DANN on the central idea of domain invariance, we focus on DANN and its adaptation to CLIP in our analysis. Notably, despite the prevalence of such DG methods, the direct application for CLIP is not well-established and remains underexplored. Naively enforcing domain invariance on foundation models like CLIP, with large pretrained knowledge, risks catastrophic forgetting.
Robust Finetuning of CLIP. The introduction of CLIP marked a significant shift in DG research. The original study radford2021learning_CLIP demonstrated impressive zero-shot classification performance across diverse benchmarks, including OOD datasets. The authors attributed this capability to CLIP learning representations that are less reliant on spurious correlations specific to downstream target datasets, as CLIP was not trained on these specific datasets during its initial pretraining.
The assumption of the inherent OOD robustness in CLIP motivated numerous methods aimed at finetuning CLIP for downstream tasks while enhancing its perceived robustness. A common approach is parameter-efficient finetuning (PEFT) strategies. An early influential study, CoOp zhou2022learning_coop , introduced learnable textual prompts, motivated by observations that manually crafted prompt ensembles improved CLIP’s zero-shot accuracy. Building on this, CoCoOp zhou2022conditional_cocoop made these prompts dynamic by conditioning them on individual image features through a cross-attention mechanism. Similarly, CLIP-Adapter gao2024clipadapter proposed adding lightweight, learnable MLP layers (adapters) to the CLIP encoders, finetuning only these small adapters instead of the entire network. Many more subsequent PEFT methods have also been explored cho2023promptstyler ; chi2024adapting ; lee2025domain ; addepalli2024leveraging ; bai2024soft ; li2022learning ; khattak2025learning ; cheng2024disentangled ; lafon2024gallop .
End-to-end finetuning methods have also been explored, yet many still depend on the original pretrained CLIP weights for regularization or guidance. Wise-FT wortsman2022robust_wise , motivated by observing that standard finetuning often degraded zero-shot OOD performance, ensembles the weights of the finetuned model with the original CLIP weights. CLIP-OOD shu2023clipood used a beta-moving average of the weights during finetuning alongside a regularization term to enhance semantic relationships learned during pretraining. MIRO cha2022domain used mutual information regularization between the finetuning model and the frozen pretrained CLIP model to retain pretrained features.
While many other methods show strong performance on OOD benchmarks, this overview highlights representative approaches, their trends, and assumptions in robust CLIP finetuning. Our work, however, questions whether current evaluation protocols are sufficiently challenging, and suggests the reliance on the pretrained weights may be suboptimal for true OOD generalization, a concern supported by evidence of domain contamination during pretraining mayilvahanan2024search . Consequently, we explore more challenging evaluations and alternative strategies for training CLIP grounded in DG principles.
3 CLIP-DCA: Disentangling Classification from Enhanced Domain-Aware Representations
The previous sections have highlighted the complexity of evaluating domain generalization in foundational models like CLIP. In light of this, we propose a novel finetuning approach, CLIP-DCA (Disentangling Classification from enhanced domain Aware representations), designed to improve robustness by balancing the trade-off between domain invariance and knowledge retention.
Figure 2: CLIP-DCA applies different sets of losses to source data images and diffusion images. For source images, accurate classification is encouraged through the classification loss between class head and text encoder ($C_{1}$). Invariance is encouraged through the disentanglement between domain and class heads ($C_{2}$). With diffusion images, domain invariance is encouraged through the disentanglement between the domain and class heads ($C_{3}$), and disentanglement between class head and text encoder ($C_{4}$). Domain awareness is encouraged through the agreement between the domain head and the text encoder ($C_{5}$), and the agreement between the text encoder and the MLLM hidden states ($C_{6}$). During inference, only the class head and text projector are used for classification.
3.1 Encouraging domain awareness and invariance simultaneously
Our key hypothesis is that domain invariance at the decision-making stage is beneficial for generalizing to unseen domains. At the same time, domain awareness is required for retaining the vast pretrained knowledge of CLIP. We achieve them simultaneously by encouraging domain awareness in the encoders, while enforcing domain invariance only in the classifier of CLIP through disentanglement. The intuition is that if a model understands what constitutes as domain-specific features, then it can learn to disregard it appropriately during classification on unseen domains.
Enforcing domain invariance in the encoder through conventional domain adversarial learning, for instance, can be harmful. Our experiments show that applying invariance directly leads to worse performance compared to standard finetuning (Figure 8). Forcing the entire model to become domain-invariant can lead to the forgetting of valuable, fine-grained features learned during the pretraining on a large dataset. Conversely, existing CLIP robust finetuning methods discourage divergence from the original pretrained model, and rely on the assumption that CLIP is inherently robust to OOD data. This assumption is challenged by our results (Figure 9) and evidence for domain contamination mayilvahanan2024search .
Figure 3: Standard CLIP inference pipeline using a dot product between image and text embeddings for classification.
Instead, we focus on enforcing domain invariance only at the final classification layer, while simultaneously encouraging the image encoder to become domain-aware. Our intuition is that a comprehensive understanding of various domains enables the model to more effectively disregard domain-specific influences during inference. The diverse set of generated diffusion images and their descriptions (detailed in Section 3.2) provides the necessary signals for enhancing this domain awareness.
To implement this, we introduce an architectural addition to the CLIP image encoder. We add an additional linear projection head, termed the image domain head ($I_{D}$), which has the same dimensionality as the original image projection head, referred to as the image class head ($I_{C}$), as shown in Figure 2. We do not add a corresponding domain head to the text encoder for two reasons. First, in most downstream classification datasets, only class names are available as text inputs, without domain descriptions. Second, textual information inherently allows for easier separation of domain and class attributes. For instance, a prompt like "a sketch of a dog" clearly distinguishes class ("dog") from domain ("sketch"). Note that for inference, the standard pipeline is used as shown in Figure 3. The domain head and other losses are not used.
During training, we use two distinct loss functions for the two types of data we use - the source dataset and generated diffusion images. We use $\ell_{a}$ to refer to agreement loss (the standard CLIP contrastive loss radford2021learning_CLIP or finetuning goyal2023finetune_flyp ). We use $\ell_{d}$ to refer to disentanglement, which we define as the squared sum of the diagonal of the cross-correlation matrix — $\ell_{d}=\sum_{i}((XY^{\mathsf{T}})_{ii})^{2}$, where $X$ and $Y$ are normalized matrices of size [batch, feature]. We also use $P_{T}$ to refer to the projected embeddings of the text encoder, and $P_{H}$ as the projected hidden states of the MLLM.
We simultaneously encourage accurate classification, domain awareness in both text and image encoders, and domain invariance at the classification stage with the following loss terms:
  1. 1.
    For the source dataset images (e.g., ImageNet, with only class labels):
    • A classification loss (i.e., the standard CLIP contrastive loss goyal2023finetune_flyp ) between the output of the image class head and the text embedding of the class name, $C_{1}:=\ell_{a}(I_{C},P_{T})$.
    • A disentanglement loss between the class and domain heads, $C_{2}:=\ell_{d}(I_{C},I_{D})$.
    • For source dataset images, we minimize the loss function $\mathcal{L}_{source}=C_{1}+C_{2}$.
  2. 2.
    For the diffusion images and their MLLM-generated style descriptions:
    • A disentanglement loss between the class head and domain head, $C_{3}:=\ell_{d}(I_{C},I_{D})$.
    • A disentanglement loss between the text embedding of style descriptions and the image class head to further encourage the class head to learn domain invariance, $C_{4}:=\ell_{d}(P_{T},I_{C})$.
    • An agreement loss between the output of the image domain head and the text embedding of the style description, enhancing domain head’s domain awareness, $C_{5}:=\ell_{a}(P_{T},I_{D})$.
    • An agreement loss between the text embedding and the corresponding projected MLLM hidden state, enhancing the text encoder’s domain awareness, $C_{6}:=\ell_{a}(P_{T},P_{H})$.
    • For diffusion images, we minimize the loss function $\mathcal{L}_{diffusion}=C_{3}+C_{4}+C_{5}+C_{6}$.
Figure 4: Pipeline for generating synthetic domain images and descriptions.
3.2 Generating diverse domains
Traditional DG benchmarks provide multi-domain datasets, enabling the learning of domain invariance. However, our evaluation setup, which involves finetuning on a single source dataset like ImageNet, lacks explicit multiple source domains, especially as the boundary for different domains becomes more vague for DG in-the-wild. Additionally, we hypothesize that to understand what constitutes as domain-specific features, a diverse number of domains are required.
To address this, we construct a small dataset with a diverse number of domains. As illustrated in Figure 4, we prompt a Multimodal Large Language Model (MLLM), specifically LLaVA liu2023visual_llava , to generate ideas of 512 distinct styles for images (e.g. "pixel art"). A text-to-image diffusion model (Stable Diffusion 3 esser2024scaling ) then generates images from these stylistic prompts. We intentionally omit any class labels during image generation to ensure the styles are not biased towards specific classes. We generate 8 images per style, creating a dataset of 4096 images. Finally, the same MLLM generates textual domain descriptions (captions) for each style. We also store the hidden state representations from the MLLM that were used to generate these style descriptions, as these will be used to encourage domain awareness in the text encoder.
4 Experimental Setup
4.1 Evaluating DG in-the-wild performance
Figure 5: PCA visualization of domains from different domain generalization datasets
While standard domain generalization benchmarks such as DigitsDG zhou2020digitsDG , PACS li2017pacs , Office-Home venkateswara2017officehome , Terra Incognita beery2018terra , FMOW-Wilds and Camelyon-Wilds koh2021fmowwilds , and ImageNet variants russakovsky2015imagenetv1 ; recht2019imagenetv2 ; hendrycks2021imagenetR ; hendrycks2021imagenetA ; wang2019imagenetSketch are widely used, we observe that the different domains within a single benchmark dataset often exhibit greater similarity to each other than to conceptually similar domains across different datasets. To quantify this, we use Spectral-normalized Neural Gaussian Process (SNGP) liu2020simple_sngp to compute the pairwise OOD scores between domains across these benchmarks. We then visualized the pairwise OOD scores with PCA. This analysis reveals significant clustering within individual benchmarks as shown in Figure 5. The clustering within benchmarks, combined with the impressively high zeroshot accuracy, and the success of transductive methods on certain DG datasets wallingford2023neuralpriming ; martin2024transductive , provides strong evidence that the current DG evaluation is not very difficult for CLIP, possibly due to domain contamination.
Consequently, we evaluate using a more challenging cross-dataset setup aiming to simulate DG in-the-wild. We finetune the model on ImageNet-1K deng2009imagenet , and evaluate its generalization capabilities across a wide range of 33 target datasets, including the standard DG benchmarks listed above. A cross-dataset evaluation is significantly more challenging compared to traditional DG setups, as it involves larger visual distribution shifts and also shifts in class labels. This evaluation also aligns with the methodologies of prior studies investigating robust CLIP finetuning zhou2022learning_coop ; zhou2022conditional_cocoop ; gao2024clipadapter ; shu2023clipood , while adding a broader coverage of domains. We use the CLIP ViT-B/32 model for all experiments. Other training details are included in the appendix.
Figure 6: OOD score of 33 target datasets against ImageNet and classification accuracy improvement over zeroshot
4.2 Measuring OODness of the target datasets
Given that our DG in-the-wild evaluation includes many target datasets with varying degrees of OODness compared to ImageNet, establishing a quantitative OOD metric is beneficial for a more holistic assessment of OOD robustness. A unique consideration for CLIP is its dual-encoder architecture. To provide a comprehensive score, we utilize OOD measures for both the image and text modalities. For the image encoder, we use SNGP liu2020simple_sngp calibrated on the ImageNet validation data to compute an OOD score for all 33 target datasets. In addition, we use a text-based OOD measure fort2021exploring_clip_text to measure OODness of class labels. This involves calculating classification probabilities on a combined label set of target dataset class names and ImageNet class names using the target domain image embeddings. The text OOD score is the summed probability assigned to the target-specific class names.
We verify that our OOD score shows a strong negative correlation (r=-0.756, p<0.001) with performance on target datasets after finetuning, as shown in Figure 6. Notably, we find that averaging the image and text OOD scores is important for accurately predicting post-finetuning accuracy. Relying solely on the image OOD score (r=-0.099) or the text OOD score (r=-0.608) yields weaker correlations, providing evidence that OOD scores in both modalities are necessary for a comprehensive understanding of OOD challenges in the context of CLIP.
4.3 Approximating unseen data through unlearning
Table 1: Performance of ZS (zero-shot), FT (Regular finetuning on GCC), and Unlearn (Regular finetuning on GCC + unlearning on DomainNet).
Metric/Data ZS FT Unlearn Imagenet IN 1 $54.2$ $52.0$ $48.8$ IN 2 $48.4$ $45.5$ $41.8$ IN Sketch $32.3$ $31.5$ $30.7$ IN A $26.2$ $19.0$ $18.2$ IN R $59.7$ $56.8$ $52.7$ DomainNet Clipart $64.3$ $67.0$ $53.0$ Infograph $41.6$ $41.0$ $34.0$ Painting $54.4$ $53.9$ $47.0$ Real $80.5$ $80.7$ $73.3$ Sketch $57.9$ $57.2$ $45.5$ Quickdraw $12.1$ $8.2$ $0.3$ Avg. on 33 $51.1$ $49.7$ $45.5$
Ideally, evaluating true DG performance would involve using a CLIP model trained without specific target-like domains. Unfortunately, retraining CLIP from scratch while omitting certain data is computationally expensive due to the scale of original training data (millions to billions of images). Public weights for selectively trained models are unavailable.
Given these constraints, we explore an approximate approach inspired by the concept of unlearning to mitigate potential domain contamination. Specifically, we adapt the adversarial learning-based unlearning method sepahvandselective for domain forgetting. We finetune CLIP goyal2023finetune_flyp using a dual objective. First, to retain general knowledge, we train on a 595,000-image subset of the CC3M dataset sharma2018conceptual_cc3m , referred to as GCC, previously used in LLaVA pretraining liu2023visual_llava , serving as a manageable proxy for CLIP’s original training data. Second, to approximate a scenario where domains similar to DomainNet are removed, we apply domain adversarial training ganin2016domain_dann on the DomainNet dataset, which we exclude from our target datasets. We attach a binary classifier to the penultimate layer of the image encoder. During training batches, this classifier is fed representations of random noise (assigned label 0) and images from DomainNet (assigned label 1). The gradient reversal layer ganin2016domain_dann forces the image encoder to learn representations that confuse this classifier, making embeddings of DomainNet images and random noise indistinguishable, thereby encouraging the model to unlearn domain-specific features from DomainNet. The unlearning occurs concurrently with standard training on the GCC dataset to preserve CLIP’s core capabilities.
We intentionally avoid unlearning the target datasets used in our evaluation. This ensures a fairer comparison against existing robust CLIP finetuning methods, many of which discourage heavily diverging from the original pretrained weights. Directly unlearning the target datasets would give the baselines an unfair disadvantage. Instead, unlearning DomainNet serves as a proxy for reducing domain contamination effects. We find that performance on DomainNet and some other datasets drops, as shown in Table 1, while retaining much of the performance on many other datasets.
5 Results and Discussion
5.1 Finetuning original pretrained CLIP
Figure 7: Performance comparison of CLIP-DCA against regular finetuning. Best-fit lines, determined by linear regression, illustrate performance trends.
We first evaluate CLIP-DCA in the context of our domain generalization in-the-wild setup, using the original pretrained CLIP weights as the starting point. As shown in Figure 7, CLIP-DCA consistently improves performance over standard finetuning across target datasets. Importantly, the best-fit line for CLIP-DCA shows a flatter slope, indicating that it is more robust to more severe OOD data compared to regular finetuning. This observation aligns with our hypothesis that encouraging domain invariance at the decision-making layer, while simultaneously encouraging domain awareness within the encoders, is crucial for robust classification on unseen distributions.
Figure 8: Comparison against more baselines
Figure 8 provides a broader comparison against additional baselines. We observe that conventional domain adversarial learning (DANN ganin2016domain_dann ), is harmful for CLIP, showing inferior performance compared to regular finetuning. This shows the potential disadvantage of enforcing domain invariance across the entire image encoder, which can lead to excessive forgetting of features learned during pretraining. This suggests the importance of approaches such as our proposed learning of targeted invariance through disentanglement.
Interestingly, on the most extremely OOD datasets, parameter-efficient finetuning (PEFT) techniques like CoOp zhou2022learning_coop and CLIP-Adapter gao2024clipadapter perform best. PEFT methods minimally change a small subset of the original CLIP weights. Consequently, their performance shows much lower variance across the datasets, with improvements (around 1-2%). It is important to note that on extreme OOD datasets, all end-to-end finetuning methods exhibit lower performance than the zero-shot CLIP baseline. While CLIP-DCA mitigates this performance drop compared to standard finetuning, it does not entirely overcome it.
This strong zero-shot performance has often been attributed to CLIP’s inherent OOD generalization capability. However, the study by mayilvahanan2024search challenges this assumption and shows that this generalization could be attributed to domain contamination. They show that when CLIP is retrained solely on natural images, its OOD performance drops to similar levels as models trained exclusively on ImageNet. This drop could offer a plausible explanation for observations like those motivating Wise-FT wortsman2022robust_wise , where standard finetuning was found to degrade OOD performance.
Table 2: Accuracy on ImageNet variants
Method V1 V2 Sketch A R
Zeroshot (unlearned) $48.8$ $41.8$ $30.7$ 18.2 $52.7$
Regular Finetune $69.8$ $58.4$ $34.7$ $15.0$ $52.6$
DANN $70.0$ $58.2$ $33.2$ $16.5$ $52.0$
CLIP Adapter $52.9$ $45.7$ $28.4$ $15.0$ $51.6$
CoOp $53.3$ $46.2$ $29.1$ $16.1$ 52.8
Wise-FT $72.9$ $61.3$ 40.0 $9.4$ $43.0$
MIRO 74.1 62.7 $35.7$ $7.3$ $33.2$
CLIP-OOD $69.0$ $58.2$ $35.3$ $15.0$ $45.8$
CLIP-DCA (Ours) 75.1 63.9 42.2 22.9 62.2
Figure 9: Comparison against baselines after unlearning
5.2 Finetuning after unlearning
To further investigate the impact of potential domain contamination and to establish a more rigorously "unseen" evaluation, we applied the unlearning procedure detailed in Section 4.3 to the pretrained CLIP model. We then finetuned this "unlearned" model on ImageNet-1K and evaluated its performance. Table 2 shows the accuracies on the ImageNet variant datasets. For this analysis, we also include several end-to-end robust finetuning methods that add a linear classifier to CLIP. Due to their architecture, these specific baselines are evaluated only on the ImageNet variants as they cannot be adapted to datasets with different class labels.
Our results show that robust end-to-end finetuning methods remain effective for datasets that are less OOD even after unlearning. For instance, MIRO cha2022domain and Wise-FT wortsman2022robust_wise outperform regular finetuning on ImageNet-V1, ImageNet-V2, and ImageNet-Sketch. However, consistent with the trends seen with the non-unlearned model, performance significantly drops on datasets with larger OOD scores, such as ImageNet-A and ImageNet-R. Similarly, PEFT methods show slight improvements over the unlearned zero-shot baseline on ImageNet-V1, V2, and Sketch, but their performance drops on ImageNet-A and R.
Figure 9 shows that the performance of all methods, even PEFT methods, further drops as OODness increases across target datasets when finetuning the unlearned model. If the unlearning process successfully reduced the knowledge of target-like domains, existing robust finetuning methods, which rely on the pretrained weights, would struggle on genuinely OOD data. These results suggest that our unlearning approach was effective in simulating a less contaminated starting point.
Table 3: Ablations on GCC inclusion. Accuracy on ImageNet variants (V1, V2, Sketch, A, R) and Avg. accuracy on 33 datasets.
Setting V1 V2 Sketch A R Avg. Zeroshot $46.0$ $40.4$ $27.4$ $15.1$ $51.6$ $45.5$ ImageNet only FLYP $69.8$ $58.4$ $34.7$ $15.0$ $52.6$ $43.6$ DANN $70.0$ $58.2$ $33.2$ $16.5$ $52.0$ $42.5$ CLIP-DCA 75.3 64.1 40.3 22.3 60.3 48.6 ImageNet+GCC FLYP $70.6$ $59.7$ $38.5$ $17.6$ $57.5$ $49.0$ DANN $70.5$ $59.4$ $38.6$ $17.4$ $57.2$ $47.5$ CLIP-DCA 75.1 63.9 42.2 22.9 62.2 52.1
With the unlearned model, CLIP-DCA shows high performance. For datasets with moderate OOD scores relative to ImageNet, CLIP-DCA achieves larger performance improvements compared to other methods. More importantly, on the extremely OOD datasets, the performance of our method remains close to the zero-shot model, without significant performance drops. This suggests that our mechanism of encouraging domain awareness while selectively enforcing invariance at the decision layer is particularly beneficial when starting from a model with reduced prior exposure to target-like domains.
5.3 Ablations
Including GCC data. When finetuning CLIP-DCA, we also use the GCC dataset – the dataset with 595,000 image-caption pairs used to prevent CLIP from collapsing during the unlearning procedure (Sec. 4.3). While the dataset is smaller than ImageNet-1K, it serves as a manageable proxy for the data CLIP was originally pretrained on. The image-caption pairs provide valuable supervision particularly for training the text encoder and possibly preventing catastrophic forgetting during finetuning on a classification datasets like ImageNet.
We study the contribution of the GCC data as shown in Table 3. A key observation is that the inclusion of GCC provides a notable benefit even for standard finetuning (FLYP) goyal2023finetune_flyp . This shows the general benefit of incorporating diverse, captioned data during finetuning. Given these benefits, an alternative or complementary approach could involve using MLLMs to generate rich textual descriptions for classes or images within the primary source dataset, similar to strategies explored in pratt2023does ; maniparambil2023enhancing , which use an LLM to describe class names. Despite the general improvements, our method consistently shows higher performance even when the GCC dataset was not included.
Table 4: Ablation of CLIP-DCA components: Domain descriptions (Domain), Disentanglement (Disent.), MLLM Hidden States (MLLM HS), and Avg. accuracy on 33 datasets.
Method / Config. Domain Disent. MLLM HS Avg. MLLM (LLaVA) - - - $24.2$ FLYP X X X $49.0$ Ours O X X $49.1$ O O X $50.8$ O X O $49.0$ Our full O O O 52.1
Different components of CLIP-DCA. We study the effect of the different components of CLIP-DCA, as shown in Table 4. We isolate the use of domain descriptions from diffusion images to train the image domain head, the disentanglement loss between the class and domain heads to encourage invariance at the classifier, and the use of MLLM hidden states to encourage domain awareness in the text encoder. Simply introducing domain descriptions to make the image encoder aware of styles, without enforcing disentanglement at the classifier, shows only a marginal improvement over the FLYP baseline, suggesting that domain awareness alone is insufficient without a mechanism to disentangle classification from it, as CLIP may otherwise struggle to disregard domain-specific features irrelevant to classification. When we incorporate the disentanglement loss to encourage domain invariance at the decision-making layer, even without explicit domain awareness in the text encoder, performance slightly improves. This is further evidence for our core hypothesis that enabling the model to disregard domain-specific features during classification is important. Attempting to make both encoders domain-aware without the disentanglement loss results in no improvement over the baseline, indicating that awareness without a mechanism for invariance can be ineffective for OOD data.
Limitations. One concern might be the reliance on synthetically generated diffusion images and MLLM-extracted features for domain awareness. However, this is mitigated by: (1) the small size of the diffusion dataset (4096 samples), (2) images synthesized using generic, class-agnostic style prompts, and (3) the MLLM processing multiple style-consistent images, which focuses it on style over objects. Furthermore, DANN ganin2016domain_dann and our ablations without disentanglement (Table 4), even with such data, fails to improve CLIP’s OOD performance (Table 3).
The role of the MLLM may also be questionable, as LLaVA internally uses a CLIP-L encoder. However, LLaVA’s poor zero-shot image classification performance (Table 4), a known issue attributed to MLLMs’ improper alignment for classification zhang2024visually , justifies not using it as a direct classifier. Instead, we use an MLLM because CLIP captures global information from images, which prioritizes overall style tong2024eyes , making its representations suitable for domain-level information. The MLLM, with its language capabilities, is then able to explain the perceived domain styles into textual descriptions and provide informative hidden state representations.
Lastly, our unlearning strategy involves making DomainNet images and random noise indistinguishable, differing from the standard approach sepahvandselective where samples are typically mapped to known OOD data. This adaptation was necessary as CLIP’s extensive web-scale pretraining makes finding truly unseen data challenging. Future work could explore more sophisticated unlearning methods for DG in-the-wild evaluation. Nevertheless, the significant degradation observed in zero-shot performance post-unlearning, and the fact that PEFT methods showed improvements on less OOD data but poorer performance on more OOD data, is evidence that our unlearning procedure functioned as intended.
6 Conclusion
In this work, we highlighted the potential limitations of current DG evaluation settings for foundation models like CLIP, which may not adequately test unseen data scenarios. We instead used a more challenging and comprehensive evaluation to simulate DG in-the-wild, with quantified OOD scores for target datasets, and an unlearning approach to further simulate unseen data. To address the challenges of DG in-the-wild, we introduced CLIP-DCA. Our method disentangles classification from domain-aware representations, motivated by the idea that while domain invariance is important for performance on unseen data, domain awareness is important to retain the vast pretrained knowledge of CLIP. Overall, our method significantly improves OOD robustness over existing baselines.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and/or introduction clearly states the claims made, including the contributions made in the paper and important assumptions and limitations.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations mentioned in separate section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: No proofs included.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All information is disclosed.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Data and code will be public.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All information is provided.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Statistical significance is provided for some results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Hardware information is mentioned.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Only publicly available datasets and models were used.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Societal impact not mentioned.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Only public generators were used.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Cited all datasets and models.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: All code and data will be made public.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No human subjects involved.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No human subjects involved.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: An opensource MLLM (multi-modal LLM) is used for our method.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.